OpenStudy (anonymous):

In a right triangle ABC, AB, the side opposite the 30 degrees angle measures 12 inches. Which is the length of the BC, the side opposite the 90 degrees angle? Answer 24 12 24square root of 2

OpenStudy (anonymous):

side opposite 90 degrees=hypotenuse (let it be x) sin(30)=12/x then x=12/sin(30)= \[x=12/(1/2)=24